Calculate the length of a solid cylinder of diameter 2.3cm and volume 435cmcube?
Answer by Hawksfan(61)

V = pi * r^2 * h
What do we know?
V = 435 cm^3
pi = 3.14
r = diameter/2 = 2.3/2 = 1.15cm
h = V/(pi * r^2) = 435 cm^3/(3.14 *1.15^2)
h = length = 104.75 cm